# Debrup Das Work done during internship at Rakuten India Enterprise Private Limited IIT Kharagpur

### v2

@@ 5 Effects of Adding Modules over LLMs
Here, we present results and analyze the impact of adding individual modules on top of the original LLM CoT variant (termed SG[Uncaptioned image]): 1) KR[Uncaptioned image] in § 5.1, 2) BS[Uncaptioned image] in §5.2, 3) PG[Uncaptioned image] in §5.4, and 4) WA[Uncaptioned image] in §5.3. For each module, we also provide ablations over different LLMs (as applicable).
@@ 5.1 LLM-Based Knowledge Retrieval (KR)
Recently, Chameleon Lu et al. (2023a) demonstrated an accuracy boost for knowledge intensive QA datasets, such as ScienceQA and TabMWP by using the KR[Uncaptioned image] module. Skills-In-Context prompting Chen et al. (2023a) also shows similar results by utilizing some basic skills (such as mathematical theorems, during generation). Following the literature, we investigate the impact of adding relevant knowledge (such as mathematical concepts and formulae) using an LLM-based KR[Uncaptioned image] module in the context of SG[Uncaptioned image] module, and examine the efficacy of the KR + SG ([Uncaptioned image] + [Uncaptioned image]) setting on the MATH dataset (Table 4). We also ablate over different LLMs (Table 2) to power the KR[Uncaptioned image] module, while fixing the SG[Uncaptioned image] module to gpt-3.5-turbo ([Uncaptioned image]).
@@ Results.
As shown in of Table 4, the extra knowledge retrieved by the KR[Uncaptioned image] module is useful only for problems in Algebra, PreAlgebra, and Probability domains. Moreover, the overall accuracy drops steadily as we change KR[Uncaptioned image]’s LLM from gpt-3.5-turbo ([Uncaptioned image]) to other variants (shown in Table 2). This indicates that, generic LLMs (such as those mentioned in Table 2) are not equipped with mathematical concepts of other domains (Precalculus, Gemetry, Number Theory, Intermediate Algebra). After analyzing different LLM variants for the KR[Uncaptioned image] module, we find that the knowledge retrieved by weaker LLMs heavily degrades performance of the downstream SG[Uncaptioned image] module. This motivated us to explore the impact of search engine-based knowledge retrieval (detailed in §5.2).
@@
Table 2: Performance of different backbone models used for KR[Uncaptioned image]  module in the KR + SG ([Uncaptioned image] + [Uncaptioned image]) setting. For all settings, we use gpt-3.5-turbo ([Uncaptioned image]) as the default LLM for the SG[Uncaptioned image] module.
5.2 Query Generation for Bing Web Search (BS)
We investigate the advantages of adding a search engine-based knowledge retrieval module (BS[Uncaptioned image]) as an alternative of KR[Uncaptioned image]  for similar questions search and concepts search before applying SG[Uncaptioned image].
@@
In Table 3, we observe that BS + SG ([Uncaptioned image] + [Uncaptioned image]) setting is a clear winner over the SG[Uncaptioned image] setting, when gpt-3.5-turbo ([Uncaptioned image]) is used for generating the BING-Web-Search-API ([Uncaptioned image]) query and getting final solution from SG[Uncaptioned image]. This holds true even if the stand-alone SG[Uncaptioned image] is varied between text-davinci-003 ([Uncaptioned image])  ($+22.5\%$) and gpt-3.5-turbo ([Uncaptioned image])  ($+4.2\%$). Thus, augmenting LLMs with knowledge (relevant to a mathematical question) retrieved from the web proves to be beneficial in improving problem solving capabilities. The use of text-davinci-003 ([Uncaptioned image]) alone or in combination with gpt-3.5-turbo ([Uncaptioned image]) for BS[Uncaptioned image] and SG[Uncaptioned image] modules, diminishes the performance of both BS + SG ([Uncaptioned image] + [Uncaptioned image]) and SG ([Uncaptioned image]) settings , which is expected Ye et al. (2023).
LLMs BS+SG WA+SG SG
([Uncaptioned image] + [Uncaptioned image]) ([Uncaptioned image] + [Uncaptioned image]) ([Uncaptioned image])
@@
${}^{\prime}$[[Uncaptioned image]] + SG ([Uncaptioned image] + [Uncaptioned image]) 55.4 23.5 58.0 22.9 32.7 42.2 17.9 39.6
@@
Table 4: Comparison of our Modular Settings to Published Baselines on MATH. We use gpt-3.5-turbo ([Uncaptioned image]) as the default LLM for each setting (except one row). For ${}^{\prime}$[[Uncaptioned image]] + SG ([Uncaptioned image] + [Uncaptioned image]) setting, we use Phind-CodeLlama-34B-V2 as the underlying LLM for the PG[Uncaptioned image] module (while keeping gpt-3.5-turbo ([Uncaptioned image]) as the default LLM for SG[Uncaptioned image] module); Alg: Algebra, P.Cal: Precalculus, P.Alg: Prealgebra, Geom: Geometry, Prob: Probability, N.Th: Number Theory, Int.Alg: Intermediate Algebra; We have taken the first four baseline results from SKiC Chen et al. (2023a), and following two baselines from Zhou et al. (2023).
@@ 5.3 Wolfram Alpha Search (WA)
We compare the performance of WA + SG ([Uncaptioned image] + [Uncaptioned image]) and SG[Uncaptioned image] settings on the MATH dataset in Table 3. We perform ablations with text-davinci-003 ([Uncaptioned image]) and gpt-3.5-turbo ([Uncaptioned image]) as the LLMs used in WA[Uncaptioned image] for query generation and answer extraction.
@@ Results.
From Table 3, we observe that WA + SG ([Uncaptioned image] + [Uncaptioned image]) outperforms the SG[Uncaptioned image] approach by 8.1%, when both WA[Uncaptioned image] and SG[Uncaptioned image] are powered by gpt-3.5-turbo ([Uncaptioned image]). This shows a clear and significant contribution of complementary strengths coming from the knowledge retrieved through Wolfram Alpha ([Uncaptioned image]). Furthermore, it is notable that the observed benefits of the WA[Uncaptioned image] module cannot be solely attributed to the characteristics of the LLMs employed for query generation or answer extraction. This is evident from the substantial performance gains (around 10.8%) achieved, even after enabling both WA[Uncaptioned image] and SG[Uncaptioned image] with a comparatively weaker model, such as text-davinci-003 ([Uncaptioned image]). Additionally, the mix of text-davinci-003 ([Uncaptioned image]) and gpt-3.5-turbo ([Uncaptioned image]) for the WA + SG ([Uncaptioned image] + [Uncaptioned image]) setting demonstrates superior performance compared to SG[Uncaptioned image] with gpt-3.5-turbo ([Uncaptioned image]), achieving improvements of 1.1% and 3.3%, respectively. Thus, showcasing meaningful positive impact of augmenting WA[Uncaptioned image] with the stand-alone SG[Uncaptioned image] module.
@@ 5.4 Python Generator (PG)
In this section, we investigate the effectiveness of the Python Generator (PG[Uncaptioned image]) module in using python code, and an interpreter to solve mathematical problems (utilizing external symbolic libraries from sympy). Following, PAL (Program Aided Language Models) Gao et al. (2023), Program of thought Chen et al. (2022), our PG[Uncaptioned image] module consists of a a program generator and an executor. The generated code and corresponding output are added in context of the next module in sequence. We present the results of the PG + SG ([Uncaptioned image] + [Uncaptioned image]) setting in Table  4 for the MATH dataset. For MATH, we present three variations: (i) PG + SG ([Uncaptioned image] + [Uncaptioned image]) with no code refinement, (ii) PG + CR+ SG ([Uncaptioned image] +[Uncaptioned image]+ [Uncaptioned image]) with code refinement, (iii) ${}^{\prime}$[[Uncaptioned image]] +SG ([Uncaptioned image] + [Uncaptioned image]) with Phind-CodeLLama-34B-V2 model used as the LLM for PG Module. We choose Phind-CodeLLama-34B-V2 for our ablation since it is the best model from the huggingface Code-LLM leaderboards. The Phind family of models are finetuned versions of CodeLlama-34B on a Phind dataset consisting of 80k high quality programming problems and solutions.
Results.
In Table 4, we observe that the PG + SG ([Uncaptioned image] + [Uncaptioned image]) setting using the sympy library without code refinement can improve upon the performance accuracy of SG [Uncaptioned image] on the MATH dataset by a margin of 10.1%. We find that a majority of problems in MATH require complex computations such as solving equations, representation of complex mathematical objects such as vectors, solving problems in Geometry, some of which are hurdles for the Solution generator module since text representations alone fail to capture such complexities. Libraries such as sympy, on the other hand, has support for symbolically representing such objects using well defined functions, classes, methods, and sub-packages. We find that this helps PG outperform SG on all mathematical types in MATH. The outcomes of our experiment with PG + CR + SG ([Uncaptioned image] + [Uncaptioned image] + [Uncaptioned image]) setting only yields marginal enhancements on overall accuracy. We also observe a drop in the accuracy by 5% when using Phind-CodeLLama-34B-V2 as the LLM in PG module.
@@ 5.5 Results of Multiple Module Experiments
We experiment with various module combinations on four datasets MATH, AQUA-RAT, GSM-8K, and MMLU-Math and report in Tabs. 4 & 6. Our findings reveal that distinct modules exhibit specialized efficacy in addressing specific categories of mathematical problems. On the MATH dataset, (1) WA[Uncaptioned image] emerges as a valuable resource for tackling intricate mathematical subdomains, particularly in Intermediate Algebra (Int.Alg) and Number Theory (N.Th). The PG+WA+SG ([Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]) setting outperforms SG([Uncaptioned image]) by 19% on Int.Alg. We conduct a qualitative analysis of PG+SG ([Uncaptioned image]+[Uncaptioned image]) on 106 randomly sampled questions from MATH spanning all types and difficulty levels, presented in Table 13. We find that the majority of errors in Int.Alg arise from Python code execution errors, clearly showcasing the inability of python code to represent complex math objects in this domain. In contrast, the WA ([Uncaptioned image]) module effectively interacts with the API using symbolic Wolfram code language to address these issues, resulting in substantial enhancements.(2) For Algebra-related problems (Prealgebra and Algebra) having complex computations, the generation of Python code guided by PG[Uncaptioned image] and the sympy library proves to be an effective choice. The WA+PG+SG ([Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]) setting elevates the performance of SG ([Uncaptioned image]) by 15.8% on Algebra. The PG+SG ([Uncaptioned image]+[Uncaptioned image]) setting performance is also significantly better compared to SG ([Uncaptioned image]) (10.4%) on Prealgebra showing the utility of code representations over natural language in this subdomain. (3) Table 8 presents an examination of the variations in accuracy among various settings as a function of the problem levels (1-5) in the MATH dataset. Our analysis reveals a consistent improvement of over 10% across all levels with diverse modular configurations. This reaffirms the importance of judiciously selecting tools and configurations based on the specific features and attributes of the given problem.
@@
Effectiveness of MathSensei on MMLU-Math.    Results in Table 6 reveal that the BS+PG+SG ([Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]) configuration enhances the accuracy of the SG ([Uncaptioned image]) setting by 3.3%. As the performance is gain is low, we further perform a type wise analysis in Table 5. We observe that, other than Formal Logic (FL), adding different modules show substantial improvements in different types, such as $17\%$ in College Math, $11.7\%$ in High School Math, $7.5\%$ in Elementary Math. More specifically we find that: (1) The PG+WA+SG ([Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]) setting improves the accuracy of the SG ([Uncaptioned image]) setting from 84.6% to 92.1% on Elementary mathematics problems. (2) Interestingly, problems in formal logic are best solved using SG ([Uncaptioned image]) alone. The drop in performance for the PG+SG ([Uncaptioned image]+[Uncaptioned image]) setting (53.9-> 49.5) is due to the inability of PG to adequately represent predicate logic, First Order Logic (FOL) sentences through python code, (3) For College Mathematics, the Wolfram Alpha ([Uncaptioned image]) module demonstrates highest efficacy, as evidenced by the substantial benefits observed in both the WA+SG ([Uncaptioned image]+[Uncaptioned image]) and WA+PG+SG([Uncaptioned image]+[Uncaptioned image]+[Uncaptioned image]) settings. Notably, WA+SG ([Uncaptioned image]+[Uncaptioned image]) outperforms the SG setting by a significant margin of 17%. Our analysis in MMLU-Math further supports the complimentary benefit of the tools used in MathSensei framework for various mathematical types.
@@
Decreased Effectiveness of MathSensei on GSM-8k, and AQUA-RAT.   From Table 6, we observe marginal improvement of using multiple modules on AQUA-RAT and GSM-8k, over the standalone SG[Uncaptioned image]  module. Both datasets mainly comprise simpler algebraic and arithmetic word problems. GSM-8K consists of problems requiring simple arithmetic operations, and mostly does not require external mathematical knowledge. The complexity in GSM-8K stems more from linguistic diversity. We conduct a case study on a randomly sampled set of 20 examples from GSM-8K, where PG + SG ([Uncaptioned image] + [Uncaptioned image]) is incorrect and SG[Uncaptioned image] is correct, we find that 18 (out of 20) have incorrect outputs generated by PG[Uncaptioned image] (due to reasoning errors)(Table 15). For all these 18 examples, the LLM generated python code tries to solve a simple problem by using complex objects in sympy, which in turn degrades the performance. For the remaining two examples, one has an execution error, while for the other one, SG[Uncaptioned image] alters the correct answer to incorrect.
@@
Similar to GSM-8K, AQUA-RAT primarily focuses on problems that require generic language-based reasoning skills. We find that settings with WA[Uncaptioned image] mostly hurt the performance compared to SG[Uncaptioned image]. This is attributed to the fact that WA[Uncaptioned image] and BS[Uncaptioned image] are unnecessary for addressing straightforward problems, and invoking them often introduces noisy and irrelevant information into the context of SG[Uncaptioned image]. As we saw previously in case of GSM-8K, a significant proportion of errors in PG + SG ([Uncaptioned image] + [Uncaptioned image]) can be linked to the application of sympy for simple problems (Table 15). These outcomes highlight the diminishing utility of employing additional modules for tasks requiring minimal external knowledge.

### v3

@@ 5 Effects of Adding Modules over LLMs
Here, we present results and analyze the impact of adding individual modules on top of the original LLM CoT variant (termed SG): KR in §5.1, BS in §5.2, PG in §5.4, and WA in §5.3. For each module, we also provide ablations over different LLMs (as applicable).
@@ 5.1 LLM-Based Knowledge Retrieval (KR)
Recently, Chameleon Lu et al. (2023a) demonstrated an accuracy boost for knowledge intensive QA datasets, such as ScienceQA and TabMWP by using the KR module. Skills-In-Context prompting Chen et al. (2023a) also shows similar results by utilizing some basic skills (such as mathematical theorems) during generation. Following the literature, we investigate the impact of adding relevant knowledge (such as mathematical concepts and formulae) using an LLM-based KR module in the context of SG module, and examine the efficacy of the KR+SG setting on the MATH dataset (Table 4). We also ablate over different LLMs (Table 2) to power the KR module, while fixing the SG module to gpt-3.5-turbo.
@@ Results.
As shown in of Table 4, the extra knowledge retrieved by the KR module is useful only for problems in Algebra, PreAlgebra, and Probability domains. Moreover, the overall accuracy drops steadily as we change KR’s LLM from gpt-3.5-turbo to other variants (shown in Table 2). This indicates that, generic LLMs (such as those mentioned in Table 2) are not equipped with mathematical concepts of other domains (Precalculus, Geometry, Number Theory, Intermediate Algebra). After analyzing different LLM variants for the KR module, we find that the knowledge retrieved by weaker LLMs heavily degrades performance of the downstream SG module. This motivated us to explore the impact of search engine-based knowledge retrieval (detailed in §5.2).
@@
Table 2: Performance of different backbone models used for KR module in the KR+SG setting. For all settings, we use gpt-3.5-turbo as the default LLM for the SG module.
5.2 Bing Web Search (BS)
We investigate the advantages of adding a search engine-based knowledge retrieval module (BS) as an alternative of KR for similar questions search and concepts search before applying SG.
@@
In Table 3, we observe that BS+SG setting is a clear winner over the SG setting, when gpt-3.5-turbo is used for generating the Bing-Web-Search-API query and getting final solution from SG. This holds true even if the stand-alone SG is varied between text-davinci-003 ($+22.5\%$) and gpt-3.5-turbo ($+4.2\%$). Thus, augmenting LLMs with knowledge (relevant to a mathematical question) retrieved from the web proves to be beneficial in improving problem solving capabilities. The use of text-davinci-003 alone or in combination with gpt-3.5-turbo for BS and SG modules, diminishes the performance of both BS+SG and SG settings, which is expected Ye et al. (2023).
Setting
LLMs ([Uncaptioned image]+[Uncaptioned image]) ([Uncaptioned image]+[Uncaptioned image]) ([Uncaptioned image])
@@
PG[[Uncaptioned image]]+SG ([Uncaptioned image]+[Uncaptioned image]) 55.4 23.5 58.0 22.9 32.7 42.2 17.9 39.6
@@
Table 4: Comparison of our Modular Settings to Published Baselines on MATH. We use gpt-3.5-turbo ([Uncaptioned image]) as the default LLM for each setting (except one row). For PG[[Uncaptioned image]]+SG ([Uncaptioned image]+[Uncaptioned image]) setting, we use Phind-CodeLlama-34B-V2 as the underlying LLM for the PG[Uncaptioned image] module (while keeping gpt-3.5-turbo ([Uncaptioned image]) as the default LLM for SG[Uncaptioned image] module); Alg: Algebra, P.Cal: Precalculus, P.Alg: Prealgebra, Geom: Geometry, Prob: Probability, N.Th: Number Theory, Int.Alg: Intermediate Algebra; We have taken the first four baseline results from SKiC Chen et al. (2023a), and following two baselines from Zhou et al. (2024).
@@ 5.3 WolframAlpha Search (WA)
We compare the performance of WA+SG and SG settings on the MATH dataset in Table 3. We perform ablations with text-davinci-003 and gpt-3.5-turbo as the LLMs used in WA for query generation and answer extraction.
@@ Results.
From Table 3, we observe that WA+SG outperforms the SG approach by 8.1%, when both WA and SG are powered by gpt-3.5-turbo. This shows a clear and significant contribution of complementary strengths coming from the knowledge retrieved through WolframAlpha. Furthermore, it is notable that the observed benefits of the WA module cannot be solely attributed to the characteristics of the LLMs employed for query generation or answer extraction. This is evident from the substantial performance gains (around 10.8%) achieved, even after enabling both WA and SG with a comparatively weaker model, such as text-davinci-003. Additionally, the mix of text-davinci-003 and gpt-3.5-turbo for the WA+SG setting demonstrates superior performance compared to SG with gpt-3.5-turbo, achieving improvements of 1.1% and 3.3%, respectively. Thus, showcasing meaningful positive impact of augmenting WA with the stand-alone SG module.
@@ 5.4 Python Generator (PG)
In this section, we investigate the effectiveness of the Python Generator (PG) module in using python code, and an interpreter to solve mathematical problems (utilizing external symbolic libraries from Sympy). Following, PAL (Program Aided Language Models) Gao et al. (2023), Program of thought Chen et al. (2023c), our PG module consists of a a program generator and an executor. The generated code and corresponding output are added in context of the next module in sequence. We present the results of the PG+SG setting in Table  4 for the MATH dataset. For MATH, we present three variations: (i) PG+SG with no code refinement, (ii) PG+CR+SG with code refinement, and (iii) PG[[Uncaptioned image]]+SG (where PG[[Uncaptioned image]] denotes the use of Phind-CodeLLama-34B-V2 model for PG. We choose Phind-CodeLLama-34B-V2 for our ablation since it is the best model from the huggingface Code-LLM leaderboards. The Phind family of models are finetuned versions of CodeLlama-34B on a Phind dataset consisting of 80k high quality programming problems and solutions.
Results
In Table 4, we observe that the PG+SG setting using the Sympy library without code refinement can improve upon the performance accuracy of SG on the MATH dataset by a margin of 10.1%. We find that a majority of problems in MATH require complex computations such as solving equations, representation of complex mathematical objects such as vectors, solving problems in Geometry, some of which are hurdles for the Solution generator module since text representations alone fail to capture such complexities. Libraries such as Sympy, on the other hand, has support for symbolically representing such objects using well defined functions, classes, methods, and sub-packages. We find that this helps PG outperform SG on all mathematical types in MATH. The outcomes of our experiment with PG+CR+SG setting only yields marginal enhancements on overall accuracy. We also observe a drop in the accuracy by 5% when using Phind-CodeLLama-34B-V2 as the LLM in PG module.
@@ 5.5 Results of Multiple Module Experiments
We experiment with various module combinations on four datasets MATH, AQUA-RAT, GSM-8K, and MMLU-Math and report in Tabs. 4 & 6. Our findings reveal that distinct modules exhibit specialized efficacy in addressing specific categories of mathematical problems. On the MATH dataset, (1) WA emerges as a valuable resource for tackling intricate mathematical subdomains, particularly in Intermediate Algebra (Int.Alg) and Number Theory (N.Th). The PG+WA+SG setting outperforms SG by 19% on Int.Alg. We conduct a qualitative analysis of PG+SG on 106 randomly sampled questions from MATH spanning all types and difficulty levels, presented in Table A. We find that the majority of errors in Int.Alg arise from python code execution errors and the inability of python code to represent complex math objects in this subdomain. In contrast, the WA module effectively interacts with the API using both natural language and symbolic queries (Table 15) to address these issues, resulting in substantial enhancements.(2) For Algebra-related problems (Prealgebra and Algebra) having complex computations, the generation of Python code guided by PG and the Sympy library proves to be an effective choice. The WA+PG+SG setting elevates the performance of SG by 15.8% on Algebra. The PG+SG setting performance is also significantly better compared to SG (10.4%) on Prealgebra showing the utility of code representations over natural language in this subdomain. (3) Table 9 presents an examination of the variations in accuracy among various settings as a function of the problem levels (1-5) in the MATH dataset. Our analysis reveals a consistent improvement of over 10% across all levels with diverse modular configurations. This reaffirms the importance of judiciously selecting tools and configurations based on the specific features and attributes of the given problem.
@@
Effectiveness of MathSensei on MMLU-Math.    Results in Table 6 reveal that the BS+PG+SG configuration enhances the accuracy of the SG setting by 3.3%. As the performance is gain is low, we further perform a type wise analysis in Table 5. We observe that, other than Formal Logic (FL), adding different modules show substantial improvements in different types, such as $17\%$ in College Math, $11.7\%$ in High School Math, $7.5\%$ in Elementary Math. More specifically we find that: (1) The PG+WA+SG setting improves the accuracy of the SG setting from 84.6% to 92.1% on Elementary Mathematics problems. (2) Interestingly, problems in Formal Logic are best solved using SG alone. The drop in performance for the PG+SG setting (53.9 -> 49.5) is due to the inability of PG to adequately represent predicate logic, First Order Logic (FOL) sentences through python code, (3) For College Mathematics, the WolframAlpha module demonstrates highest efficacy, as evidenced by the substantial benefits observed in both the WA+SG and WA+PG+SG settings. Notably, WA+SG outperforms the SG setting by a significant margin of 17%. Our analysis in MMLU-Math further supports the complementary benefit of the tools used in MathSensei framework for various mathematical types.
@@
Decreased Effectiveness of MathSensei on GSM-8K, and AQUA-RAT.   From Table 6, we observe marginal improvements of using multiple modules on AQUA-RAT and GSM-8K, over the standalone SG module. Both datasets comprise simpler algebraic and arithmetic word problems. GSM-8K consists of problems requiring simple arithmetic operations such as addition, subtraction, etc. and its complexity stems from linguistic diversity. We conduct a case study on a randomly sampled set of 20 examples from GSM-8K, where PG+SG is incorrect and SG is correct, we find that 18 (out of 20) have incorrect outputs generated by PG (due to reasoning errors) (Table A).
[…]
SG ([Uncaptioned image]): PG+SG ([Uncaptioned image]+[Uncaptioned image]) BS+SG ([Uncaptioned image]+[Uncaptioned image]) WA+SG ([Uncaptioned image]+[Uncaptioned image])
[…]
SG module misses the repeated characters in NINE.
from sympy import *
word = ’NINE’
numperm=factorial(len(word)); print(numperm);
[…]
PG module ignores the repeated characters.
Retrieved Example: In the word "BANANA", there are 6 letters but only 3 distinct letters …repeated letters.
[…]
BS module retrieves similar examples and relevant formula from the Web.
[…]
WA module generates the correct query: Permutations[NINE] to solve the given question and final SG answer is grounded on WA output.
Table 7: Qualitative Analysis of the Responses generated by different Settings for a given MATH example.
@@
For all these 18 examples, the LLM generated python code tries to solve a simple problem by using complex objects in Sympy, which in turn degrades the performance. For the remaining two examples, one has an execution error, while for the other one, SG alters the correct PG answer to incorrect. Similar to GSM-8K, AQUA-RAT primarily focuses on problems that require generic language-based reasoning skills. We find that settings with tools mostly hurt the performance compared to SG. This is attributed to the fact that WA and BS are unnecessary for addressing straightforward problems, and invoking them often introduces noisy and irrelevant information into the context of SG. As we saw previously in case of GSM-8K, a significant proportion of errors in PG+SG ([Uncaptioned image]+[Uncaptioned image]) can be linked to the application of Sympy for simple problems (Table A). These outcomes highlight the diminishing utility of employing additional modules for tasks requiring minimal external knowledge.
5.6 Insights from Qualitative Analysis of Modules
We consider an example from the MATH dataset and present a qualitative analysis of the responses generated by different settings in Table 7. We observe that SG and PG+SG are unable to capture the fine-grained nuances in the input question (repetition of characters in the word "NINE"), leading to reasoning errors. On the other hand, the BS+SG and WA+SG settings avoid committing such errors. This demands a need for a careful examination of the strengths and limitations of the individual modules, which we discuss in details in this section.
Bing Web Search (BS).
Previous investigations of retrieval-augmented generation (RAG) Lewis et al. (2021) and Self-RAG Asai et al. (2023) have shown that conditional generation using retrieval-based approaches improves factuality in knowledge intensive tasks such as question answering, fact verification, etc. We observe similar benefits of employing retrieval-based methods in the domain of complex mathematical reasoning. The BS module retrieves useful information (such as formulas, concepts and similar questions) from the Web, improving the effectiveness of the downstream SG module. As shown in Table 7, the BS module retrieves an informative example of a similar question (permutations of the letters in the word "BANANA", having repeated characters) and the correct formula of permutations with repetitions, which aids the SG module in correctly reasoning about the final solution. However, our current implementation of the BS module also has certain limitations. We directly use the raw output returned by the Bing Web Search API v7, which is noisy in certain cases. Additionally, we do not employ any critique mechanism to check the relative importance of multiple pieces of the retrieved information. We also observe a significant reduction in performance on GSM-8K after adding the BS module to SG. This calls for a component which can effectively decide when it is required to retrieve knowledge and when it is not necessary (future research).
WolframAlpha (WA).
The WA module overcomes the limitations of SG by harnessing the computational power and intelligence of the WolframAlpha engine. In cases, where the query to the WolframAlpha-API is syntactically and logically correct for solving a mathematical question, the returned answer is guaranteed to be correct, which is then processed by the SG module to compile the final answer. From Figure 6, we observe maximum benefit of WolframAlpha module for problems in the subdomains of Algebra and Intermediate Algebra (primarily for difficulty levels greater than one). We demonstrate the utility and limitations of the WA module in Tables 16 and A, respectively. The limitations of the WA module are mostly associated with: (1) Logical errors in LLM-generated WA API Queries (Example 1 in Table A), (2) Wrong interpretation of correct WA response by the downstream SG module (Example 2 in Table A), (3) Single line WA response, which restricts the ability of downstream SG module to generate step-by-step reasoning (Example 3 in Table A).
Python Generator (PG).
As mentioned in Section 5.4, the Sympy library offers strong capabilities to the PG module for MATH. The deterministic program executor also helps in avoiding common errors committed by the standalone SG module. We found the PG module to be most useful in solving Algebra, Prealgebra and Number Theory problems. The example presented in Table 17 demonstrates the advantage of using the PG module over the SG module. The primary errors of the PG module in Intermediate Algebra and Prealgebra are mainly due to inability of the generated python code to express complex objects (syntax errors) and boundary cases, respectively. In the case of Geometry and Precalculus problems, a large proportion of errors are caused due to lack of understanding the plots/figures (expressed in latex format) accompanying the question. Table A presents examples of common syntactical errors (from the MATH dataset) of the PG module, and Table A summarizes the different error types of PG+SG setting. Similar to PoT Chen et al. (2023c), we found the PG module to be less effective for simpler arithmetic problems (removal of Sympy improves the performance by 2-3%). However, the overall performance of SG and PG+SG still remains quite similar. In Table 18, we present an example from GSM-8K where PG+SG commits a reasoning error while the SG setting is correct.